A string in the shape of a 2 X 8 inch rectangle can be shaped into a square whose side is
A) 4 inches B) 5 inches C) 16 inches D) 20 inches

Answer by checkley77(12569)

2*2+2*8=4+16=20 inches is the perimeter of the rectangle.
A square has 4 equal sides.
20/4=5 inches for each side of the square.